If 1.80 kg glucose C6H12O6 is burned completely, how many kg of CO2 is produced?






* Skill - Calculate the amount of water produced in this case.






1
Expert's answer
2022-01-28T10:33:03-0500

Mass of water =0.084


Mass=mole × mm


Mass = 0.084×18 = 1.512


Mass of H2O = 1.512kg